# Prompting in Practice: Investigating Software Developers’ Use of Generative AI Tools

### v1

@@ 3.1.3. Changes in Development Approaches with GenAI
When asked if GenAI tools had changed their approach to software development, the majority of respondents (55/71, 77.5%) indicated that they had experienced at least some change, while 7.0% were unsure and 15.5% reported no change. We asked respondents who reported experiencing changes to their development approach to elaborate on those changes. Nine participants reported using GenAI tools for tasks that previously required internet searches (e.g., finding library documentation or searching Stack Overflow for code examples), and four reported using GenAI to assist in information retrieval tasks within their code base. Six indicated that GenAI can provide a starting point for projects, such as helping with ideation (3), initial project design (5), or writing boilerplate code (6). Five respondents noted that GenAI allowed them to shift development focus from repetitive coding tasks to higher-level tasks such as architecture design, and four respondents claimed that GenAI tools allowed them to write more documentation. Still others indicated that GenAI helps produce cleaner, more organized code (3), reduces the focus and memory required to complete coding tasks (2), allows for development in unfamiliar programming languages (2), and can assist in project configuration (2).
GenAI tool adoption has created new processes in developer workflows. For example, six respondents spend more time reviewing and verifying AI-produced code. Four focus more on writing good prompts, while three have incorporated interactive prompting processes into their workflows.
Finding 5: Developers reported that GenAI has transformed their development workflows in diverse ways, ranging from replacing traditional search practices to supporting project design, reducing repetitive effort, and broadening programming capabilities, ultimately enabling a shift toward higher-level, more strategic software engineering tasks. In addition, participants noted that GenAI introduced new practices in their development workflow, such as validating AI-generated outputs and employing various prompt engineering strategies.
@@ 3.2. RQ2: Prompting and Conversation Strategies
This section examines the specific prompting techniques and conversation strategies that software developers use when interacting with GenAI tools. We first analyze participants’ familiarity with prompting strategies, then investigate what contextual information developers include when prompting GenAI for different SE tasks, what conversation strategies they use, and how developers react to inadequate output from GenAI tools.
@@ 3.2.1. Prompting Technique Familiarity
Participants were asked about three prompting techniques identified by White et al. (Meta Language Creation, Output Automator, and Persona(White et al., 2023) in addition to Meta-Prompting and Few-Shot Learning (Schulhoff et al., 2025). We also included the Output Style and Condition Check techniques, which we did not find adequately represented in existing work. Definitions for these techniques were provided to participants and can be found in Table 3.
Prompting Technique Technique Description
Condition Check Produce certain type(s) of output when specific condition(s) are met
Few-Shot Learning Providing examples to guide generation
Meta Language Creation Adjusts/Adds to the semantics of words, phrases, or symbols
Meta-Prompting Having AI suggest better prompts for specific tasks
Output Automator Generating scripts to implement solutions
Output Style Make the AI’s output follow a particular format or style
Persona Ask AI to complete a task while acting as a certain role or character
@@
The Output Style (54%) and Persona (44%) strategies were most likely to be used regularly, among participants who used them. Despite being least popular overall, 42% of respondents who used the Meta Language Creation strategy do so regularly. Respondents most often reported using the Output Automator (75%), Condition Check (68%), and Meta-Prompting (68%) strategies only occasionally.
@@
We found that none of respondents’ self-reported GenAI proficiency, years of development experience, or GenAI proficiency and experience impacted relative familiarity ratings of prompting techniques (e.g., across proficiency groups, Output Style always had a higher familiarity than Meta Language Creation).
Table 4 reports the percentage of respondents who are familiar with, and use, the seven prompting techniques across GenAI proficiency and SE experience categories. Technique usage rates are further reported by occasional and regular use. The table reports total percentages by aggregating response counts across the techniques. The results show a clear pattern: both technique familiarity and usage (including occasional and regular use) consistently increase with self-reported proficiency. In other words, participants who considered themselves more proficient were also more likely to use prompting techniques more frequently.
In contrast, analysis by developer experience reveals a different trend. With the exception of veterans, all experience groups show similar familiarity and usage rates. Veteran are the least familiar with prompting techniques, but among those who are familiar, they are the most likely to use them. Analyzing occasional versus regular technique usage, developers across all experience levels show consistent rates, except for intermediate developers, who reported over 60% increase in regular usage compared with their peers. This suggests that intermediate developers may be at a career stage where they are both open to adapting their workflows and sufficiently experienced to integrate GenAI effectively.
Finding 7: Familiarity and usage of prompting techniques increase with GenAI proficiency rather than years of experience. Veteran developers are the least familiar with prompting techniques, yet the most likely to use the ones they know. While most experience groups use prompting at similar rates, intermediate developers stand out as especially active adopters, suggesting they are mid-career professionals that are uniquely positioned to integrate GenAI into their workflows.
@@ 3.2.2. Prompt Information Inclusion
We describe the prompt information developers use for each of the six SE tasks we studied.
@@
When prompting a GenAI tool for code generation, 76% of respondents (66) indicated that they often include example inputs and expected outputs. Respondents also indicated including instructions for which libraries the model should use (53%), a description of the existing codebase (45%), implementation alternatives to either consider or avoid (41%), error handling expectations (39%), coding style guidelines (32%), deployment constraints (32%), and performance/optimization requirements (27%).
@@
When using GenAI tools for refactoring tasks, respondents (23) most often included the original code with explanatory comments (78%), a description of the existing issues or refactoring goals (65%), and examples of similar code that they found to be well structured (61%). Others opted to include architectural constraints and style guidelines (48%), a specific refactoring pattern to apply (48%), or code that should remain unchanged during the refactoring (43%). The least common information to be supplied in respondent prompts included unit tests (22%) and performance requirements (17%).
@@
Of the 28 respondents that used GenAI for testing tasks, 82% prompted the model with the code to be tested along with specifications. Respondents also included example test cases (46%), testing preferences (43%), information on edge cases or expected behaviors (39%), existing test suite examples (39%), and mocking instructions (36%) in their prompts. Least likely to be included were test coverage goals (25%), environmental requirements (25%), and performance or resource constraints (14%).
@@
When using GenAI for debugging tasks, nearly all respondents (94%) provide the model with full error messages or stack traces. Similarly, 81% often included related logs or console output. Less common context given to the model included environmental details (50%), previously attempted solutions (44%), steps-to-reproduce (41%), system architecture information (31%), and visual evidence of issue manifestation (e.g., screenshots) (25%). Only two respondents (6%) included version control history.
@@
For documentation tasks, respondents (34) reported that they often provided the model with the code snippet along with a functionality explanation (71%). Some also provided examples highlighting the desired documentation style (53%), usage examples/scenarios (41%), context about the target audience (38%), existing documentation in need of updating (38%), format requirements (35%), guidance on domain-specific terminology (26%), or information on project standards and conventions (15%).
@@
Lastly, when using GenAI for code review, respondents (23) predominantly provided the model with the code to be review with some additional context (78%). Beyond this, they included project coding standards (43%), a description of the intended feature or purpose (35%), or performance expectations (30%), security/compliance requirements (22%), warnings about common pitfalls in the domain or codebase (22%), the code of related components (17%), or feedback from previous reviews (17%).
@@ 3.2.3. Conversation Strategy Adoption
We examined how developers structured their conversations with GenAI tools. Participants rated the use frequency of nine conversation strategies we defined (seen in Table 5) on a four-point Likert scale from “Never” to “Often.” This analysis reveals which strategies survey developers prefer.
@@
Table 5. Conversation strategy descriptions
Figure 4 reveals that the most frequently used strategy was Incremental refinement (incrementally refining the same prompt) followed closely by Feedback loop (using potentially different prompts to achieve the desired result). These strategies reflect developers’ preference for multi-turn conversations where they can guide and refine GenAI responses progressively. However, Single comprehensive prompt is the third most popular and does not reflect this iterative workflow. The least popular strategies were Comparative analysis, Template-based approach, and Exploratory dialogue. Despite their lower adoption rates, these strategies still showed meaningful usage patterns, with at least 40% of developers employing them “Sometimes” or “Often”.
@@ 3.2.4. Error Handling Strategies
We asked participants to indicate which error handling strategies they use when they deem GenAI responses as inadequate. Participants were provided with nine strategies and could select none or any number of strategies, including none. Developers converge overwhelmingly on two complementary strategies (Figure 5): providing additional context and pointing out specific issues, with 79% of frequent users employing both approaches. Figure 5 shows that reformulating requests, adding constraints, including example solutions, breaking down complex problems, and providing feedback on partial solutions show more moderate adoption, while requesting alternatives or explanations remain far less popular. However, the 13 respondents who often use this least popular strategy are likely to use every other strategy: none chose “Never” in response to any option. There is a similar trend for the 15 respondents who frequently request alternatives: only three selected “Never” for requesting explanations and not for any other strategy.
Finding 8: Developers who reported utilizing less popular error handling strategies typically also used all other provided strategies at least rarely. This suggests that strategies are used as complements to others.
The remainder of this section reveals connections between conversation structuring and error handling approaches, focusing exclusively on users who responded “Often” to both types of strategies. 80% of developers who frequently use Incremental refinement also provide additional context when errors occur. Similarly, 83% of users who provide feedback on partial solutions frequently utilize Feedback loop. Those who use step-by-step guidance often break down complex problems when handling errors (50%). These correlations show consistent developer strategies, which could inform future tool design: for example, a tool might activate additional context provision settings if a developer is aware of their preference for an iterative strategy, or adjust a prompting agent’s writing style depending on the user.
Two distinct conversation preferences emerge: Comprehensive users who rely on single detailed prompts (29/72) versus Iterative users who employ multi-turn refinement (48/72), with an overlap of 19 respondents.
Finding 9: Developers consistently reported combining related conversation structuring and error handling strategies, demonstrating preferences for two distinct GenAI interaction styles. Comprehensive prompters try to include as much detail in prompts as possible, while Iterative prompters start conversations planning to add additional context as needed while adapting to LLM outputs, with the Iterative style being significantly more popular.
@@ 3.2.5. Exchange Count
Finding 10: No developers indicated that they could consistently complete tasks with GenAI in a single exchange, while extended conversations with GenAI tools correlate with higher productivity and perceived skill.
@@ 3.3. RQ3: GenAI Reliability and Common Issues
This section examines which SE subtasks developers perform using GenAI tools, their perceptions of these tools’ reliability across SE scenarios, and how often they encounter various issues. Our results highlight where GenAI tools excel and struggle.
@@ 3.3.1. Task Engagement Patterns
Finding 11: For code generation, documentation, and testing, respondents generally reported using GenAI for implementation-focused (lower-level) sub-tasks rather than for design and analysis (higher-level) sub-tasks.
@@
3.3.2. GenAI Perceived Reliability
To assess developers’ trust in GenAI tools, we analyzed reported reliability ratings for 12 scenarios spanning all six studied SE tasks. Participants rated each scenario on a 4-point scale from “Very unreliable” to “Very reliable,” with “Unsure/Haven’t tried” responses excluded from our analysis.
Figure 6 shows the perceived reliability of using GenAI for various tasks. Respondents reported that documentation tasks were the most reliable to complete with GenAI. In particular, maintaining documentation accuracy received the highest avg. rating (3.06, n=32), followed by documenting implementation rationale (2.77, n=31). Respondents showed moderate confidence in using GenAI for testing and debugging tasks, including test case design and coverage (2.64, n=25) and root cause analysis of existing errors (2.57, n=30). The tasks reported as least reliable by developers often involved complex analysis or reasoning. These included the assessment of non-functional requirements (2.14, n=14), the modernizing of outdated code patterns (2.32, n=22), and the implementation of complex algorithms (2.35, n=60). In particular, GenAI tools may struggle with non-functional requirements since many of these require a deep understanding of domain knowledge from a complex field, such as security or license compliance.
Finding 12: Respondents perceived GenAI to be less reliable for software tasks that required complex reasoning, non-trivial analysis, and domain knowledge.
@@ 3.3.2. GenAI Perceived Reliability
Despite nearly equivalent totals for Proficient, Very Proficient, and Maximally Proficient users, Somewhat Proficient users have the lowest at 1.98, largely due to an outlying quantity of “Very Unreliable” selections. This is expected, since it is unlikely that respondents would assign themselves low proficiency if they were able to obtain reliable AI responses. The most proficient users were unlikely to classify GenAI as “Very Unreliable” or “Very Reliable” for any task, suggesting that they have at least occasionally witnessed AI fail in optimal scenarios and succeed in sub-optimal ones.
In contrast, our analysis by experience level shows no clear patterns. The most- and least-experienced developers were more likely to find GenAI reliable than those in the middle. This may be correlated with Juniors’ and Veterans’ proportionally lower familiarity with prompting techniques, but that does not explain why Intermediate developers consider AI more reliable than Senior developers.
Finding 13: Senior developers (with 11-20 years of development experience) find GenAI least reliable on average, while veterans (with 21+ years) find it most reliable on average.

### v2

@@ 3.1.3. Changes in Development Approaches with GenAI
When asked if GenAI tools had changed their approach to software development, the majority of respondents (55/71, 77.5%) indicated that they had experienced at least some change, while 7.0% were unsure and 15.5% reported no change. We asked respondents who reported experiencing changes to their development approach to elaborate on those changes. Nine participants reported using GenAI tools for tasks that previously required internet searches (e.g., finding library documentation or searching Stack Overflow for code examples), and four reported using GenAI to assist in information retrieval tasks within their code base. Six indicated that GenAI can provide a starting point for projects, such as helping with ideation (3), initial project design (5), or writing boilerplate code (6). Five respondents noted that GenAI allowed them to shift development focus from repetitive coding tasks to “higher-level” tasks such as architecture design, and four respondents claimed that GenAI tools allowed them to write more documentation. Still others indicated that GenAI helps produce cleaner, more organized code (3), reduces the focus and memory required to complete coding tasks (2), allows for development with unfamiliar programming languages (2), and can assist in project configuration (2).
GenAI tool adoption has also created new processes in practitioner workflows. For example, six respondents spend more time reviewing and verifying AI-produced code. Four focus more on writing good prompts, while three have incorporated interactive prompting processes into their workflows.
Finding 5: Practitioners reported that GenAI has transformed their development workflows in diverse ways, ranging from replacing traditional search practices to supporting project design, reducing repetitive effort, and broadening programming capabilities, ultimately enabling a shift toward higher-level, more strategic software engineering tasks. In addition, participants noted that GenAI introduced new practices in their development workflow, such as validating AI-generated outputs and employing various prompt engineering strategies.
@@ 3.2. RQ2: Prompting and Conversation Strategies
We examine the specific prompting techniques and conversation strategies that software practitioners use when using GenAI tools.
@@ 3.2.1. Prompting Technique Familiarity
Participants were asked about three prompting techniques identified by White et al. (Meta Language Creation, Output Automator, and Persona(White et al., 2023) in addition to Meta-Prompting and Few-Shot Learning (Schulhoff et al., 2025). We included two additional techniques, Output Style and Condition Check, which we did not find adequately represented in prior work. Technique definitions were provided to participants and can be found in Table 3.
Technique Description Condition Check Produce certain output when specific condition(s) are met Few-Shot Learning Providing examples to guide generation Meta Lang. Creation Adjusts the semantics of words, phrases, or symbols Meta-Prompting Having AI suggest better prompts for specific tasks Output Automator Generating scripts to implement solutions Output Style Make the AI’s output follow a particular format or style Persona Ask AI to complete a task while acting as a certain role
@@
The Output Style (54%) and Persona (44%) strategies were most likely to be used regularly, among participants who used them. Despite being least popular overall, 42% of respondents who used Meta Language Creation, use it regularly. Respondents most often reported using the Output Automator (75%), Condition Check (68%), and Meta-Prompting (68%) strategies only occasionally.
@@
None of respondents’ self-reported GenAI proficiency, years of software development experience, or GenAI proficiency and experience impacted relative familiarity ratings of prompting techniques (e.g., across proficiency groups, Output Style always had a higher familiarity than Meta Language Creation).
Table 4 reports the proportion of respondents who are familiar with, and use, the seven prompting techniques across GenAI proficiency and SE experience categories. Technique usage rates are further reported by occasional and regular use. The table reports total percentages by aggregating response counts across the techniques. The results show a clear pattern: both technique familiarity and usage (including occasional and regular use) consistently increase with self-reported proficiency. In other words, participants who considered themselves more proficient were also more likely to use prompting techniques more frequently.
In contrast, analysis by practitioner experience reveals a different trend. With the exception of veterans, all experience groups show similar familiarity and usage rates. Veteran are the least familiar with prompting techniques, but among those who are familiar, they are the most likely to use them. Analyzing occasional versus regular technique usage, practitioners across all experience levels show consistent rates, except for intermediate practitioners, who reported over 60% increase in regular usage compared with their peers. This suggests that intermediate practitioners may be at a career stage where they are both open to adapting their workflows and sufficiently experienced to integrate GenAI effectively.
Finding 7: Familiarity and usage of prompting techniques increase with GenAI proficiency rather than years of developement experience. Veteran practitioners are the least familiar with prompting techniques, yet the most likely to use the ones they know. While most experience groups use prompting at similar rates, intermediate practitioners stand out as especially active adopters, suggesting they are mid-career professionals that are uniquely positioned to integrate GenAI into their workflows.
@@ 3.2.2. Prompt Information Inclusion
We describe the prompt information practitioners use for each of the six SE tasks we studied.
@@
When prompting a GenAI tool for code generation, 76% of respondents (66) indicated that they often include example inputs and expected outputs. Respondents also indicated including instructions for which libraries the model should use (53%), a description of the existing codebase (45%), implementation alternatives to consider or avoid (41%), error handling expectations (39%), coding style guidelines (32%), deployment constraints (32%), and performance/optimization requirements (27%).
@@
When using GenAI tools for refactoring tasks, respondents (23) most often include the original code with explanatory comments (78%), a description of the existing issues or refactoring goals (65%), and examples of similar code that they find to be well structured (61%). Others opt to include architectural constraints and style guidelines (48%), a specific refactoring pattern to apply (48%), or code that should remain unchanged during the refactoring (43%). The least common information supplied in respondent prompts include unit tests (22%) and performance requirements (17%).
@@
Of the 28 respondents who use GenAI for testing tasks, 82% prompt the model with the code to be tested along with specifications. Respondents also include example test cases (46%), testing preferences (43%), information on edge cases or expected behaviors (39%), existing test suite examples (39%), and mocking instructions (36%) in their prompts. The least common to be included are test coverage goals (25%), environmental requirements (25%), and performance or resource constraints (14%).
@@
When using GenAI for debugging tasks, nearly all respondents (94%) provide the model with full error messages or stack traces. Similarly, 81% often include related logs or console output. Less common context given to the model includes environmental details (50%), previously attempted solutions (44%), (bug) reproduction steps (41%), system architecture information (31%), and visual evidence of issue manifestation (e.g., screenshots) (25%). Only two respondents (6%) included version control history.
@@
For documentation tasks, respondents (34) reported that they often provide the model with the code snippet along with a functionality explanation (71%). Some also provide examples highlighting the desired documentation style (53%), usage examples/scenarios (41%), context about the target audience (38%), existing documentation in need of updates (38%), format requirements (35%), guidance on domain-specific terminology (26%), or information on project standards and conventions (15%).
@@
Lastly, when using GenAI for code review, respondents (23) predominantly provide the model with the code to be reviewed with some additional context (78%). Beyond this, they include project coding standards (43%), a description of the intended feature or purpose (35%), or performance expectations (30%), security/compliance requirements (22%), warnings about common pitfalls in the domain or codebase (22%), the code of related components (17%), or feedback from previous reviews (17%).
@@ 3.2.3. Conversation Strategy Adoption
We examined how practitioners structure their conversations with GenAI tools. Participants rated the use frequency of nine conversation strategies we defined (seen in Table 5) on a four-point Likert scale from “Never” to “Often.” This analysis reveals which strategies survey respondents prefer.
@@
Table 5. Conversation strategy with GenAI tools
Figure 4 reveals that the most frequently used strategy was Incremental refinement (incrementally refining the same prompt) followed closely by Feedback loop (using potentially different prompts to achieve the desired result). These strategies reflect practitioners’ preference for multi-turn conversations where they can guide and refine GenAI responses progressively. However, Single comprehensive prompt is the third most popular and does not reflect this iterative workflow. The least popular strategies were Comparative analysis, Template-based approach, and Exploratory dialogue. Despite their lower adoption rates, these strategies still showed meaningful usage patterns, with at least 40% of practitioners employing them “Sometimes” or “Often”.
@@ 3.2.4. Error Handling Strategies
We asked participants to indicate which error handling strategies they use when they deem GenAI responses as inadequate. Participants were provided with nine strategies and could select none or any number of strategies. Practitioners converge on two complementary strategies (Figure 5): providing additional context and pointing out specific issues, with 79% of frequent users employing both approaches. Figure 5 shows that reformulating requests, adding constraints, including example solutions, breaking down complex problems, and providing feedback on partial solutions show more moderate adoption, while requesting alternatives or explanations remain far less popular. However, the 13 respondents who often use this least popular strategy are likely to use every other strategy: none chose “Never” in response to any option. There is a similar trend for the 15 respondents who frequently request alternatives: only three selected “Never” for requesting explanations and not for any other strategy.
Finding 8: Practitioners who reported utilizing less popular error handling strategies typically also used all other provided strategies at least rarely. This suggests that strategies are used as complements to others.
We identified connections between conversation structuring and error handling approaches, focusing exclusively on users who responded “Often” to both types of strategies. 80% of practitioners who frequently use Incremental refinement also provide additional context when errors occur. Similarly, 83% of users who provide feedback on partial solutions frequently utilize Feedback loop. Those who use step-by-step guidance often break down complex problems when handling errors (50%). These correlations show consistent practitioner strategies, which could inform future tool design: for example, a tool might activate additional context provision settings if a practitioner is aware of their preference for an iterative strategy, or adjust a prompting agent’s writing style depending on the user.
@@ 3.2.5. Exchange Count
Finding 9: No practitioners reported consistently completing tasks with GenAI in a single exchange, while longer interactions appear associated with higher productivity and perceived proficiency.
@@ 3.3. RQ3: GenAI Reliability and Common Issues
This section examines which SE subtasks practitioners perform using GenAI tools, their perceptions of these tools’ reliability across SE scenarios, and how often they encounter various issues. Our results highlight where GenAI tools excel and struggle.
@@ 3.3.1. Task Engagement Patterns
Finding 10: For code generation, documentation, and testing, respondents generally used GenAI for implementation-focused (lower-level) sub-tasks rather than for design and analysis (higher-level) sub-tasks.
@@
3.3.2. GenAI Perceived Reliability
To assess practitioners’ trust in GenAI tools, we analyzed reported reliability ratings for 12 scenarios spanning all six studied SE tasks. Participants rated each scenario on a 4-point scale from “Very unreliable” to “Very reliable,” with “Unsure/Haven’t tried” responses excluded from our analysis.
Figure 6 shows the perceived reliability of using GenAI for various tasks. Respondents reported that documentation tasks were the most reliable to complete with GenAI. In particular, maintaining documentation accuracy received the highest avg. rating (3.06, n=32), followed by documenting implementation rationale (2.77, n=31). Respondents showed moderate confidence in using GenAI for testing and debugging tasks, including test case design and coverage (2.64, n=25) and root cause analysis of existing errors (2.57, n=30). The tasks reported as least reliable by practitioners often involved complex analysis or reasoning. These included the assessment of non-functional requirements (2.14, n=14), the modernizing of outdated code patterns (2.32, n=22), and the implementation of complex algorithms (2.35, n=60). In particular, GenAI tools may struggle with non-functional requirements since many of these require a deep understanding of domain knowledge from a complex field, such as security or license compliance.
Finding 11: Respondents perceived GenAI as less reliable for software tasks requiring complex reasoning, analysis, and domain knowledge.
@@ 3.3.2. GenAI Perceived Reliability
Despite nearly equivalent totals for Proficient, Very Proficient, and Maximally Proficient users, Somewhat Proficient users have the lowest at 1.98, largely due to an outlying quantity of “Very Unreliable” selections. This is expected, since it is unlikely that respondents would assign themselves low proficiency if they were able to obtain reliable AI responses. The most proficient users were unlikely to classify GenAI as “Very Unreliable” or “Very Reliable” for any task, suggesting that they have at least occasionally witnessed GenAI fail in optimal scenarios and succeed in sub-optimal ones.
In contrast, our analysis by experience level shows no clear patterns. The most- and least-experienced respondents were more likely to find GenAI reliable than those in the middle. This may be correlated with Juniors’ and Veterans’ proportionally lower familiarity with prompting techniques, but that does not explain why Intermediates consider AI more reliable than Seniors.
Finding 12: Senior practitioners (with 11-20 years of development experience) find GenAI least reliable on average, while veterans (with 21+ years) find it most reliable on average.